OpenStudy (anonymous):

Find the unknown number in each proportion. 9/x=15/3 Can you Please explain how u got x

OpenStudy (anonymous):

cross-multiply to get 9*3 = 15*x 27 = 15x 27/15 = x 9/5 = x Is it clear???????

OpenStudy (anonymous):

thx!